Create a 3x3 magic square using the number 3,5,7,9,11,13,15,17,and 19
Found 2 solutions by solver91311, Edwin McCravy:
Answer by solver91311(24713)

Look at your grid as a wrap around. The bottom connects to the top and the sides connect to each other.

Start in the upper center cell with the smallest number. Then move up and right to insert the next number. You can't go up because you are already at the top, but remember the wrap around feature -- up right takes you to the lower right hand cell for the next higher number. up right from there takes you to the middle row left (wrap right to left) for the third number. If ever you get to a point where the cell you want to go to is already filled, go straight down. Continue in this pattern until all cells are filled. The result will be a magic square.

This works for any N X N grid where N is odd.

Answer by Edwin McCravy(20060)

Edwin's solution:
Create a 3x3 magic square using the number 3,5,7,9,11,13,15,17,and 19


| 5 |15|13|
----------
|19 |11| 3|
----------
| 9 | 7|17|

All 3 rows add up to 33:

5+15+13=33
19+11+3=33
 9+7+17=33

All 3 columns add up to 33:

 5+19+9=33
15+11+7=33
13+3+17=33

Both diagonals add up to 33:

5+11+17=33
13+11+9=33

Edwin